# Prompting Large Language Models for Recommender Systems: A Comprehensive Framework and Empirical Analysis

### v1

@@ 2.2.1. LLM as Recommendation Model
This paradigm takes the LLM as a recommender system. Employing diverse strategies like pre-training, fine-tuning, or prompting, LLMs can combine general knowledge with input data to yield personalized recommendations for users (Hou et al., 2023; Dai et al., 2023a; Ji et al., 2023). Due to the variety of recommendation tasks, LLM as recommendation model can be categorized into two types: discriminative recommendation and generative recommendation.
@@
$\bullet$ Discriminative recommendation instructs LLMs to make recommendation decisions on the given candidate items, usually focusing on item scoring (Fu et al., 2023) and re-ranking tasks (Dai et al., 2023a). For Click-Through Rate (CTR) prediction tasks, Liu et al.  (Liu et al., 2023c) designed specific zero-shot and few-shot prompts to evaluate abilities of LLMs on rating predictions. LLMs were required to assign a score for the item according to the previous rating history of users and the score range given in prompts, while the result indicated that LLMs can outperform classical rating methods (e.g., MF and MLP) in few-shot conditions (Liu et al., 2023c). Kang et al.  (Kang et al., 2023) further formulated the rating prediction task as multi-class classification and regression task, investigating the influence of model size on recommendation performance. Different from these methods, Hou et al.  (Hou et al., 2023) structured a re-ranking task, employing in-context learning approaches for LLMs to rank items in the candidate pool. Previous studies highlighted the sensitivity of LLMs to the sequence of interaction histories provided in prompts (Ma et al., 2023), which can be alleviated by strategies such as recency-focused prompting (Hou et al., 2023).
@@
$\bullet$ Generative recommendation requires LLMs to generate items recommended to users, either from candidate item lists within prompts or from LLMs with general knowledge (Li et al., 2023l). GenRec (Ji et al., 2023) leveraged the contextual comprehension ability of LLMs to transform interaction histories into formulated prompts for next-item predictions. To address instances where GenRec might propose items absent in candidate lists, GPT4Rec (Li et al., 2023m) came up with the method that used BM25 algorithm to retrieve the most similar item in candidate item lists with the item generated by LLMs. In addition to top-n recommendations, LLMs can be leveraged for generative tasks such as explainable recommendations (Lei et al., 2023; Colas et al., 2023; Li et al., 2023f; Yang et al., 2023b; Luo et al., 2023a) and review summarization (Geng et al., 2022; Liu et al., 2023d; Wang et al., 2023c). Moreover, with the incredible abilities in dialogue comprehension and communication, LLMs are naturally considered as the backbone of conversational and interactive recommender systems. ChatRec (Gao et al., 2023) designed an interactive recommendation framework based on ChatGPT, which can comprehend requirements of users through multi-turn dialogues and traditional recommendation models. Moreover, RecLLM (Friedman et al., 2023) combined the dialogue management module with a ranker module and a controllable LLM-based user simulator to generate synthetic conversations for tuning system modules. Apart from these methods, InteRecAgent (Huang et al., 2023) employed LLMs as the brain and recommender models as tools, combining their respective strengths to create an interactive recommender system (Huang et al., 2023). As a conversational recommender system, InteRecAgent enabled traditional recommender systems to become interactive systems with a natural language interface through the integration of LLMs.
@@
There are mainly two paradigms for adapting LLMs as recommenders, i.e., non-tuning paradigm and tuning paradigm.
@@
$\bullet$ Non-tuning paradigm keeps parameters of LLMs fixed and extracts the general knowledge of LLMs with prompting strategies. Existing work of non-tuning paradigm focuses on designing appropriate prompts to stimulate recommendation abilities of LLMs (Yao et al., 2023; Wang et al., 2023c; Li et al., 2023k). Liu et al.  (Liu et al., 2023c) proposed a prompt construction framework to evaluate abilities of ChatGPT on five common recommendation tasks, each type of prompts contained zero-shot and few-shot versions. Hou et al.  (Hou et al., 2023) not only used prompts to evaluate abilities of LLMs on sequential recommendation, but also introduced recency-focused prompting and in-context learning strategies to alleviate order perception and position bias issues of LLMs. ChatRec (Gao et al., 2023) and InteRecAgent (Huang et al., 2023) mentioned above are also within the classic non-tuning paradigm.
@@
$\bullet$ Tuning paradigm aims to update parameters of LLMs to inject recommendation capabilities into LLM itself. The tuning strategies include fine-tuning (Zheng et al., 2023; Hu et al., 2021) and instruction tuning (Luo et al., 2023b; Qiu et al., 2023). P5 (Geng et al., 2022) proposed five types of instructions targeting at different recommendation tasks to fine-tune a T5 (Raffel et al., 2020) model. The instructions were formulated based on conventional recommendation datasets with designed templates, which equipped LLMs with generation abilities for unseen prompts or items (Geng et al., 2022). InstructRec (Zhang et al., 2023f) further designed abundant instructions for tuning, including 39 manually designed templates with preference, intention, task form and context of a user. Compared with these methods, TallRec (Bao et al., 2023b) used LoRA (Hu et al., 2021), a parameter-efficient tuning method, to handle the two-stage tuning for LLMs. It was first fine-tuned on general data of Alpaca (Taori et al., 2023), and then further fine-tuned with the historical information of users.
@@
Although LLMs as recommendation models present a way of utilizing the common knowledge of LLMs, it still encounters some problems to be coped with. Due to the high computational cost (Touvron et al., 2023b; Liao et al., 2023) and slow inference time (Li et al., 2023b), LLMs are struggled to be efficient enough compared to traditional recommendation methods (Hou et al., 2023; Gao et al., 2023). Additionally, constraints on input sequence length will limit the amount of external information (e.g., candidate item lists) (Lin et al., 2023b), leading to the degrading performance of LLMs in scenarios such as sequential recommendation. Furthermore, since information in recommendation tasks is challenging to be expressed in natural language (Lin et al., 2023c; Yuan et al., 2023a), it is hard to formulate appropriate prompts that make LLMs truly understand what they are required to do, leading to unexpected performance.
2.2.2. LLM Improves Recommendation Models
@@
This method mainly utilizes LLMs to generate auxiliary information to enhance the performance of recommendation models (Du et al., 2023; Wang et al., 2023a; Wei et al., 2023), based on the reasoning abilities and common knowledge. The research on how to improve recommendation models with LLMs can be divided into three categories, i.e., LLM as feature encoder, LLM for data augmentation and LLM co-optimized with domain-specific models.
@@
$\bullet$ LLM as feature encoder. The representation embeddings of users and items are important factors in classical recommender systems (Rendle et al., 2009; He et al., 2020). LLMs serving as feature encoders can generate related textual data of users and items, and enrich their representations with semantic information. U-BERT (Qiu et al., 2021) injected user representations with user review texts, item review texts and domain IDs, augmenting the contextual semantic information in user vectors. Wu et al.  (Wu et al., 2021b), on the other hand, employed language models to generate item representations for news recommendation. With the development of LLMs and prompting strategies, BDLM (Zhang et al., 2023d) constructed the prompt consisting of interaction and contextual information into LLMs, and obtained top-layer feature embeddings as user and item representations.
@@
$\bullet$ LLM for data augmentation. For this paradigm, LLMs are required to generate auxiliary textual information for data augmentation (Agrawal et al., 2023; Wei et al., 2023; Liu et al., 2023b, a). By using prompting or in-context learning strategies, the related knowledge will be extracted out in different text forms to facilitate recommendation tasks (Xi et al., 2023; Du et al., 2023; Wang et al., 2023a). One form of auxiliary textual information is summarization or text generation, enabling LLMs to enrich representations of users or items (Wang, 2023). For example, Du et al.  (Du et al., 2023) proposed a job recommendation model which utilized the capability of LLMs for summarization to extract user information and job requirements. Considering item descriptions and user reviews, KAR (Xi et al., 2023) extracted the reasoning knowledge on user preferences and the factual knowledge on items through specifically designed prompts, while SAGCN (Liu et al., 2023b) utilized a chain-based prompting strategy to generate semantic information. Another form of using the textual features generated from LLMs is for graph augmentation in the recommendation field. LLMRG (Wang et al., 2023a) leveraged LLMs to extend nodes in recommendation graphs. The resulting reasoning graph was encoded using GNN, which served as additional input to enhance sequential models. LLMRec (Wei et al., 2023) adopted three types of prompts to generate information for graph augmentation, including implicit feedback, user profile and item attributes.
@@
$\bullet$ LLM co-optimized with domain-specific models. The categories mentioned above mainly focus on the impact of common knowledge for domain-specific models (Wang, 2023). However, LLM itself often struggles to handle domain-specific tasks due to the lack of task-related information (Yao et al., 2023; Kang et al., 2023). Therefore, some studies conducted experiments to bridge the gap between LLMs and domain-specific models. BDLM (Zhang et al., 2023d) proposed an information sharing module serving as an information storage mechanism between LLMs and domain-specific models. The user embeddings and item embeddings stored in the module were updated in turn by the LLM and the domain-specific model, enhancing the performance of both sides. CoLLM (Zhang et al., 2023b) combined LLMs with a collaborative model, which formed collaborative embeddings for LLM usage. By tuning LLM and collaborative module, CoLLM showed great improvements in both warm and cold-start scenarios. In conversational recommender systems, approaches such as ChatRec (Gao et al., 2023) and InteRecAgent (Huang et al., 2023) considered LLMs as the backbone, and leveraged traditional recommendation models for candidate item retrieval.
@@
In addition to the context limitation and computational cost of LLMs (Touvron et al., 2023b), the paradigm that LLM improves recommendation models also encounters other problems. (1) Although LLMs can enhance offline recommender systems to avoid online latency, this paradigm also limits the ability of LLMs to model real-time collaborative filtering information, neglecting the key factor for recommendation (Wang, 2023; Xi et al., 2023; Wei et al., 2023). (2) Feature encoding, data augmentation, and collaborative training inevitably expose the user data to LLMs, which may bring privacy, security and ethical issues (Weidinger et al., 2021; Shen et al., 2023; Carranza et al., 2023).
@@ 2.2.3. LLM as Recommendation Simulator
Due to the gap between offline metrics and online performance of recommendation methods (Pan et al., 2022; He et al., 2017), it is necessary for the designed approach to get intents of users by simulating real-world elements. In this way, LLM as the recommendation simulator is introduced by taking LLMs as the foundational architecture of generative agents, and agents simulate the virtual users in the recommendation environment (Wang et al., 2023i; Zhang et al., 2023e, c).
Recently, there emerged a lot of work studying the performance of LLMs as the recommendation simulator. Agent4rec (Zhang et al., 2023e) was a movie simulator consisting of two core fractions: LLM-empowered generative agents and recommendation environment. The work equipped each agent with user profile, memory and actions modules, mapping basic behaviors of real-world users. AgentCF (Zhang et al., 2023c), on the other hand, considered not only users but also items as agents. It captured the two-sided relations between users and items, and optimized these agents by prompting them to reflect on and adjust the misleading simulations collaboratively (Zhang et al., 2023c). Moreover, in addition to behaviors within the recommender system, RecAgent (Wang et al., 2023i, g) took external influential factors of user agent simulation into account, such as friend chatting and social advertisement. In order to describe users accurately, RecAgent applied five features for users, and implemented two global functions including real-human playing and system intervention to operate agents flexibly.
Although LLM as recommendation simulator aims to imitate real-world recommendation behaviors to enhance the recommendation performance, it still has deficiencies in some aspects. Firstly, since current work is mainly demo systems that operate a few agents (Wang et al., 2023i, g), there still exists a gap between virtual agent environment and real-world practical recommendation applications, which requires further research and development. Additionally, LLMs may arise privacy and safety concerns. Many studies take ChatGPT as the architecture of agents, presenting security risks to the recommended information for users (Zhang et al., 2023c). Moreover, Zhang et al.  (Zhang et al., 2023e) have explored that hallucination in LLMs can exert huge impact on recommendation simulations. The LLM sometimes fails to accurately simulate human users, such as providing inconsistent score for an item and fabricating non-existent items for rating.

### v2

@@ 2.2.1. LLMs as Recommender Systems
This paradigm takes LLMs as recommender systems. Employing diverse strategies like pre-training, fine-tuning, or prompting, LLMs can combine general knowledge with input data to yield personalized recommendations for users (Hou et al., 2024; Dai et al., 2023a; Ji et al., 2023). Due to the variety of recommendation tasks, LLMs as recommender systems can be categorized into two types: discriminative recommendation and generative recommendation.
@@
$\bullet$ Discriminative recommendation instructs LLMs to make recommendation decisions on the given candidate items, usually focusing on item scoring (Fu et al., 2023) and re-ranking tasks (Dai et al., 2023a). For Click-Through Rate (CTR) prediction tasks, TALLRec (Bao et al., 2023b) used the recommendation data to fine-tune LLaMA (Touvron et al., 2023a) as the CTR prediction model, showcasing the powerful capabilities of LLMs in few-shot learning. Liu et al.  (Liu et al., 2023c) designed specific zero-shot and few-shot prompts to evaluate abilities of LLMs on rating predictions. LLMs were required to assign a score for the item according to the previous rating history of users and the score range given in prompts, while the result indicated that LLMs can outperform classical rating methods in few-shot conditions (Liu et al., 2023c). Kang et al.  (Kang et al., 2023) further formulated the rating prediction task as multi-class classification and regression task, investigating the influence of model size on recommendation performance. Different from these methods, Hou et al.  (Hou et al., 2024) structured a re-ranking task, employing ICL approaches for LLMs to rank items in the candidate pool. Previous studies highlighted the sensitivity of LLMs to the sequence of interaction histories provided in prompts (Ma et al., 2023), which can be alleviated by strategies such as recency-focused prompting (Hou et al., 2024).
@@
$\bullet$ Generative recommendation requires LLMs to generate items recommended to users, either from candidate item lists within prompts or from LLMs with general knowledge (Li et al., 2024d). GenRec (Ji et al., 2023) leveraged the contextual comprehension ability of LLMs to transform interaction histories into formulated prompts for next-item predictions. To address instances where GenRec might propose items absent in candidate lists, GPT4Rec (Li et al., 2023k) came up with the method that used Best Matching (BM25) (Robertson and Zaragoza, 2009) algorithm to retrieve the most similar item in candidate item list with the item generated by LLMs. Faced with the challenge of item identifiers for compatibility with LLMs in recommender systems, LC-Rec (Zheng et al., 2024) employed vector quantization to convert natural languages into trainable tokens, enabling the alignment with input formats of LLMs. To further address the limitations of previous methods reliant on historical data, To make the recommendation results generated by LLMs break through the limitations of existing items, SpecGR (Ding et al., 2024) introduced a draft-then-verify framework to recommend new items in inductive settings. In addition to top-n recommendations, LLMs can be leveraged for generative tasks such as explainable recommendations (Lei et al., 2023; Colas et al., 2023; Li et al., 2023f; Yang et al., 2023b; Luo et al., 2024) and review summarization (Geng et al., 2022; Liu et al., 2023d; Wang et al., 2024c). Moreover, with the incredible abilities in dialogue comprehension and communication, LLMs are naturally considered as the backbone of conversational and interactive recommender systems. ChatRec (Gao et al., 2023) designed an interactive recommendation framework based on ChatGPT, which can comprehend requirements of users through multi-turn dialogues and traditional recommendation models. Moreover, RecLLM (Friedman et al., 2023) combined the dialogue management module with a ranker module and a controllable LLM-based user simulator to generate synthetic conversations for tuning system modules. Apart from these methods, InteRecAgent (Huang et al., 2023) employed LLMs as the brain and recommender models as tools, combining their respective strengths to create an interactive recommender system (Huang et al., 2023). As a Conversational Recommender System (CRS), InteRecAgent enabled traditional recommender systems to become interactive systems with a natural language interface through the integration of LLMs.
@@
The methods for adapting LLMs as recommender systems mainly have two paradigms, i.e., non-tuning paradigm and tuning paradigm as follows.
@@
$\bullet$ Non-tuning paradigm keeps parameters of LLMs fixed and extracts the general knowledge of LLMs with prompting strategies. Existing work of non-tuning paradigm focuses on designing appropriate prompts to stimulate recommendation abilities of LLMs (Yao et al., 2023; Wang et al., 2024c; Li et al., 2023j). Liu et al.  (Liu et al., 2023c) proposed a prompt construction framework to evaluate abilities of ChatGPT on five common recommendation tasks, each type of prompts contained zero-shot and few-shot versions. Hou et al.  (Hou et al., 2024) not only used prompts to evaluate abilities of LLMs on sequential recommendation, but also introduced recency-focused prompting and ICL strategies to alleviate order perception and position bias issues of LLMs. ChatRec (Gao et al., 2023) and InteRecAgent (Huang et al., 2023) mentioned above are also within the classic non-tuning paradigm.
@@
$\bullet$ Tuning paradigm aims to update parameters of LLMs to inject recommendation capabilities into LLM itself. The tuning strategies include fine-tuning (Zheng et al., 2024; Hu et al., 2021; Bao et al., 2023a; Zhang et al., 2024b) and instruction tuning (Luo et al., 2023; Qiu et al., 2023). P5 (Geng et al., 2022) proposed five types of instructions targeting at different recommendation tasks to fine-tune a T5 (Raffel et al., 2020) model. The instructions were formulated based on conventional recommendation datasets with designed templates, which equipped LLMs with generation abilities for unseen prompts or items (Geng et al., 2022). InstructRec (Zhang et al., 2023e) further designed abundant instructions for tuning, including 39 manually designed templates with preference, intention, task form and context of a user. Compared with these methods, TallRec (Bao et al., 2023b) used Low-Rank Adaptation of LLMs (LoRA) (Hu et al., 2021), a parameter-efficient tuning method, to handle the two-stage tuning for LLMs. To further align LLMs with personalized human preferences, recent research (Chen et al., 2024c; Liao et al., 2024; Gao et al., 2024) has applied post-training techniques such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) to fine-tune LLMs for recommendation, thereby integrating negative samples and user preference information within the tuning process.
@@
Although LLMs as recommender systems present a way of utilizing the common knowledge of LLMs, it still encounters problems to be coped with. Due to the high computational cost (Touvron et al., 2023b; Liao et al., 2023) and slow inference time (Li et al., 2023c), LLMs are struggled to be efficient enough compared to traditional recommendation methods (Hou et al., 2024; Gao et al., 2023). Additionally, constraints on the input sequence length will limit the amount of external information (e.g., candidate item lists) (Lin et al., 2024a), leading to degrading performance of LLMs in scenarios such as sequential recommendation. Furthermore, since information in recommendation tasks is challenging to be expressed in natural language (Lin et al., 2023b; Yuan et al., 2023a), it is hard to formulate appropriate prompts that make LLMs truly understand what they are required to do.
2.2.2. LLMs Improve Recommender Systems
@@
This method mainly utilizes LLMs to generate auxiliary information to enhance the performance of recommendation models (Du et al., 2023; Wang et al., 2023a; Wei et al., 2024b), based on the reasoning abilities and common knowledge. The research on how to improve recommendation models with LLMs can be divided into three categories, i.e., LLMs as feature encoder, LLMs for data augmentation and LLMs co-optimized with domain-specific models.
@@
$\bullet$ LLMs as feature encoder. The representation embeddings of users and items are important factors in classical recommender systems (Rendle et al., 2009; He et al., 2020). LLMs serving as feature encoders can generate related textual data of users and items, and enrich their representations with semantic information. U-BERT (Qiu et al., 2021) injected user representations with user review texts, item review texts and domain IDs, augmenting the contextual semantic information in user vectors. Wu et al.  (Wu et al., 2021b), on the other hand, employed language models to generate item representations for news recommendation. With the development of LLMs and prompting strategies, BDLM (Zhang et al., 2023c) constructed the prompt consisting of interaction and contextual information into LLMs, and obtained top-layer feature embeddings as user and item representations, injecting user-item interaction information into embeddings.
@@
$\bullet$ LLMs for data augmentation. For this paradigm, LLMs are required to generate auxiliary textual information for data augmentation (Agrawal et al., 2023; Wei et al., 2024b; Liu et al., 2023b, a). By using prompting or ICL strategies, the related knowledge will be extracted out in different text forms to facilitate recommendation tasks (Xi et al., 2023; Du et al., 2023; Wang et al., 2023a). One form of auxiliary textual information is summarization or text generation, enabling LLMs to enrich representations of users or items (Wang, 2024). For example, Du et al.  (Du et al., 2023) proposed a job recommendation model which utilized the capability of LLMs for summarization to extract user information and job requirements. Mysore et al.  (Mysore et al., 2023), on the other hand, used the ICL strategy to instruct LLMs for narrative queries according to the interaction history of users, which was then utilized as the source data for training narrative driver recommendation models. Considering item descriptions and user reviews, KAR (Xi et al., 2023) extracted the reasoning knowledge on user preferences and the factual knowledge on items through specifically designed prompts, while SAGCN (Liu et al., 2023b) utilized a chain-based prompting strategy to generate semantic information. Another form of using the textual features generated from LLMs is for graph augmentation in the recommendation field. LLMRG (Wang et al., 2023a) leveraged LLMs to extend nodes in recommendation graphs. The resulting reasoning graph was encoded using GNN, which served as additional input to enhance sequential models. LLMRec (Wei et al., 2024b) adopted three types of prompts to generate information for graph augmentation, including implicit feedback, user profile and item attributes.
@@
$\bullet$ LLMs co-optimized with domain-specific models. The categories mentioned above mainly focus on the impact of common knowledge for domain-specific models (Wang, 2024). However, LLM itself often struggles to handle domain-specific tasks due to the lack of task-related information (Yao et al., 2023; Kang et al., 2023). Therefore, some studies conducted experiments to bridge the gap between LLMs and domain-specific models. BDLM (Zhang et al., 2023c) proposed an information sharing module serving as an information storage mechanism between LLMs and domain-specific models. The user embeddings and item embeddings stored in the module were updated in turn by the LLM and the domain-specific model, enhancing the performance of both sides. CoLLM (Zhang et al., 2023b) combined LLMs with a collaborative model, which formed collaborative embeddings for LLM usage. By tuning LLM and collaborative module, CoLLM showed great improvements in both warm and cold-start scenarios. In CRS, approaches such as ChatRec (Gao et al., 2023) and InteRecAgent (Huang et al., 2023) considered LLMs as the backbone, and leveraged traditional recommendation models for candidate item retrieval.
@@
In addition to the context limitation and computational cost of LLMs (Touvron et al., 2023b), the paradigm that LLM improves recommendation models also encounters other problems. (1) Although LLMs can enhance offline recommender systems to avoid online latency, this paradigm also limits the ability of LLMs to model real-time collaborative filtering information, neglecting the key factor for recommendation (Wang, 2024; Xi et al., 2023; Wei et al., 2024b). (2) Feature encoding, data augmentation, and collaborative training inevitably expose the user data to LLMs, which may bring privacy, security and ethical issues (Weidinger et al., 2021; Shen et al., 2023; Carranza et al., 2023).
@@ 2.2.3. LLM as Recommendation Simulator
Due to the gap between offline metrics and online performance of recommendation methods (Pan et al., 2022; He et al., 2017), it is necessary for the designed approach to get intents of users by simulating real-world elements. In this way, LLM as the recommendation simulator is introduced by taking LLMs as the foundational architecture of generative agents, and agents simulate the virtual users in the recommendation environment (Wang et al., 2023h; Zhang et al., 2023d, 2024a).
Recently, there emerged a lot of work studying the performance of LLMs as the recommendation simulator. Agent4rec (Zhang et al., 2023d) was a movie simulator consisting of two core fractions: LLM-empowered generative agents and recommendation environment. The work equipped each agent with user profile, memory and actions modules, mapping basic behaviors of real-world users. AgentCF (Zhang et al., 2024a), on the other hand, considered not only users but also items as agents. It captured the two-sided relations between users and items, and optimized these agents by prompting them to reflect on and adjust the misleading simulations collaboratively (Zhang et al., 2024a). Moreover, in addition to behaviors within the recommender system, RecAgent (Wang et al., 2023h, e) took external influential factors of user agent simulation into account, such as friend chatting and social advertisement. In order to describe users accurately, RecAgent applied five features for users, and implemented two global functions including real-human playing and system intervention to operate agents flexibly.
Although LLM as recommendation simulator aims to imitate real-world recommendation behaviors to enhance the recommendation performance, it still has deficiencies in some aspects. Firstly, since current work is mainly demo systems that operate a few agents (Wang et al., 2023h, e), there still exists a gap between virtual agent environment and real-world practical recommendation applications, which requires further research and development. Additionally, LLMs may arise privacy and safety concerns. Many studies take ChatGPT as the architecture of agents, presenting security risks to the recommended information for users (Zhang et al., 2024a). Moreover, Zhang et al.  (Zhang et al., 2023d) have explored that hallucination in LLMs can exert huge impact on recommendation simulations. The LLM sometimes fails to accurately simulate human users, such as providing inconsistent score for an item and fabricating non-existent items for rating.